OpenStudy (anonymous):

Its been a while since I did maths at this level and have an assignment asking me the following. If I could get some advice this would be great. Re-arrange Equation 1, to derive Equation 2. Show your working: V = C*M+X*(1-M) M = (V - X)/(C - X)

OpenStudy (alekos):

do you still need help on this?

OpenStudy (anonymous):

Yes please :) lol

OpenStudy (alekos):

v = cm +x(1-m) v = cm + x -xm v = m(c-x) +x v-x = m(c-x) m(c-x) = v-x m = (v-x)/(c-x)

OpenStudy (anonymous):

Thank you, most appreciated. Makes sense what you have done. Thanks

OpenStudy (alekos):

your welcome